# Cluster-cyclic condition of skew-symmetrizable matrices of rank 3 via the Markov constant

### v1

@@ 10.2. Fixed points and finite orbits
An element $M=\left(\begin{smallmatrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{smallmatrix}\right)\in\mathcal{M}^{+}$ is a fixed point for the action $\Gamma$ if and only if it holds that
@@
Assume $|\Gamma(M)|<\infty$ and set $S=\mathrm{Sk}(M)$. Then, since the map $\mathrm{Sk}:\Gamma(M)\to\Gamma(S)$ is surjective, we have $|\Gamma(S)|\leq|\Gamma(M)|<\infty$. Thus, there is a maximal element $S^{\prime}\in\Gamma(S)$. Then, since $S^{\prime}$ satisfies (M1) or (M2), $S^{\prime}\leq\gamma_{i}(S^{\prime}),\gamma_{j}(S^{\prime})$ must hold for at least two indices $i,j=1,2,3$. Without loss of generality, we may assume $i=1$ and $j=2$. Then, since $S^{\prime}$ is maximal, it also holds that $\gamma_{1}(S^{\prime}),\gamma_{2}(S^{\prime}),\gamma_{3}(S^{\prime})\leq S^{\prime}$ and, by $S^{\prime}\leq\gamma_{1}(S^{\prime}),\gamma_{2}(S^{\prime})$, we have $\gamma_{1}(S^{\prime})=\gamma_{2}(S^{\prime})=S^{\prime}$. Set $S^{\prime}=(p,q,r)$. The equalities $\gamma_{1}(S^{\prime})=\gamma_{2}(S^{\prime})=S^{\prime}$ imply $p=q$ and $r=2$. Moreover, since $\gamma_{3}(S^{\prime})=(p,p,p^{2}-2)\leq S^{\prime}$, we need $p^{2}-2p\leq 2$. Since $p\geq 2$, the inequality $p^{2}-2\leq 2$ implies $p=2$. Thus, a maximum element $M^{\prime}$ should satisfy $\mathrm{Sk}(M^{\prime})=(2,2,2)$. By Lemma 10.8, $M^{\prime}$ must be expressed as in (60). Moreover, since $\Gamma(M)=\Gamma(M^{\prime})=\{M^{\prime}\}$, $M=M^{\prime}$ holds. ∎
@@ 10.3. Surjectivity of the Markov constants
In [BBH11, Thm. 1.2 (3)], they showed that there are no integral cluster-positive elements $S$ with $40<C(S)<$ 4. Moreover, an element $S$ satisfying (M1) and $C(S)=0$ is only $S=(3,3,3)$. On the other hand, by extending integer skew-symmetrizable matrices $\mathcal{M}_{\mathbb{Z}}^{\mathrm{CP}}$ (or $\widehat{\mathcal{S}}^{\mathrm{CP}}$), we can find some elements with $40\leq C(S)<$ 4 as in Example 10.18. More strongly, we obtain the following result.
@@
Next, we give all elements $S\in\widehat{\mathcal{S}}^{\mathrm{CP}}$ satisfying $40\leq C(S)<$ 4 and (M1). The classification in this region is important by the following property.
@@
(a) There exists a real number $R>2$ satisfying $C(R,R,R)=C$.
@@
Let $S=(p,q,r)\in\mathcal{S}^{+}$ satisfy (M1) and $p\geq q\geq r$. Let $C\in\mathbb{Z}_{<4}$. If $C=C(S)$, then we have $2<r$ anf $p\leq r\sqrt{\frac{r^{2}-C}{r^{2}-4}}$.
@@
Consider an element $S=(p,q,r)\in\widehat{\mathcal{S}}^{\mathrm{CP}}$ satisfying (M1) and $40\leq C(S)<$ 4. The real number $R>2$ satisfying $C(R,R,R)=3R^{2}-R^{3}=0$ is $R=3$. Thus, by Lemma 10.15, under the assumption $p\geq q\geq r$, we need $2<r\leq 3$. Since $r^{2}\in\mathbb{Z}$, we have $\sqrt{5}\leq r\leq 3$. Moreover, by setting $C=C(S)$, we have $p\leq r\sqrt{\frac{r^{2}-C}{r^{2}-4}}\leq 3\times\sqrt{\frac{3^{2}-0}{\sqrt{5}^{2}-4}}=9$ by Lemma 10.16. In particular, by considering $p^{2},q^{2},r^{2}\in\mathbb{Z}$, the number of possible $S$ are at most $5\times 81^{2}$. (By considering other conditions such as $pqr\in\mathbb{Z}$ and $r\leq q\leq p$, the number of elements which we need to check is much smaller than $5\times 81^{2}$.) We may write all cluster-positive orbits satisfying $40\leq C(S)<$ 4 as follows:

### v2

@@ 9.2. Fixed points and finite orbits
An element $M=\left(\begin{smallmatrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{smallmatrix}\right)\in\mathcal{M}^{+}$ is a fixed point for the action $\Gamma$ if and only if $\mathrm{Sk}(M)=(2,2,2)$, that is,
@@
Assume $|\Gamma(M)|<\infty$ and set $S=\mathrm{Sk}(M)$. Then, since the map $\mathrm{Sk}:\Gamma(M)\to\Gamma(S)$ is surjective, we have $|\Gamma(S)|\leq|\Gamma(M)|<\infty$. Thus, there is a maximal element $S^{\prime}\in\Gamma(S)$. Then, since $S^{\prime}$ satisfies (M1) or (M2), $S^{\prime}\leq\gamma_{i}(S^{\prime}),\gamma_{j}(S^{\prime})$ must hold for at least two indices $i,j=1,2,3$. Without loss of generality, we may assume $i=1$ and $j=2$. Then, since $S^{\prime}$ is maximal, it also holds that $\gamma_{1}(S^{\prime}),\gamma_{2}(S^{\prime}),\gamma_{3}(S^{\prime})\leq S^{\prime}$ and, by $S^{\prime}\leq\gamma_{1}(S^{\prime}),\gamma_{2}(S^{\prime})$, we have $\gamma_{1}(S^{\prime})=\gamma_{2}(S^{\prime})=S^{\prime}$. Set $S^{\prime}=(p,q,r)$. The equalities $\gamma_{1}(S^{\prime})=\gamma_{2}(S^{\prime})=S^{\prime}$ imply $p=q$ and $r=2$. Moreover, since $\gamma_{3}(S^{\prime})=(p,p,p^{2}-2)\leq S^{\prime}$, we need $p^{2}-2\leq 2$. Since $p\geq 2$, the inequality $p^{2}-2\leq 2$ implies $p=2$. Thus, a maximum element $M^{\prime}$ should satisfy $\mathrm{Sk}(M^{\prime})=(2,2,2)$. By Lemma 9.8, $M^{\prime}$ is a fixed point and $M^{\prime}$ must be expressed as in (59). Moreover, since $\Gamma(M)=\Gamma(M^{\prime})=\{M^{\prime}\}$, $M=M^{\prime}$ holds. ∎
@@ 9.3. Surjectivity of the Markov constants
[…]
@@
Next, we give all elements $S\in\widehat{\mathcal{S}}^{\mathrm{CP}}$ satisfying $0\leq C(S)<4$ and (M1). The classification in this region is important by the following property.
@@
(a) There exists a real number $R>2$ satisfying $C(R,R,R)=C$. Moreover, this $R$ is unique.
@@
[…]
@@
Consider an element $S=(p,q,r)\in\widehat{\mathcal{S}}^{\mathrm{CP}}$ satisfying (M1) and $0\leq C(S)<4$. The real number $R>2$ satisfying $C(R,R,R)=3R^{2}-R^{3}=0$ is $R=3$. Thus, by Lemma 9.15, under the assumption $p\geq q\geq r$, we need $2<r\leq 3$. Since $r^{2}\in\mathbb{Z}$, we have $r=\sqrt{5},\sqrt{6},\sqrt{7},\sqrt{8},3$. Moreover, by setting $C=C(S)$, we have $p\leq r\sqrt{\frac{r^{2}-C}{r^{2}-4}}\leq 3\times\sqrt{\frac{3^{2}-0}{\sqrt{5}^{2}-4}}=9$ by Lemma 9.16. In particular, by considering $p^{2},q^{2},r^{2}\in\mathbb{Z}$, the number of possible $S$ are at most $5\times 81^{2}$. (By considering other conditions such as $pqr\in\mathbb{Z}$ and $r\leq q\leq p$, the number of elements which we need to check is much smaller than $5\times 81^{2}$.) We may write all cluster-positive orbits satisfying $0\leq C(S)<4$ as follows:
@@
10. The case (B)
Recall that there are two classes (A) and (B) among cluster-positive elements in $\mathcal{S}$. (See Definition 7.6.) For the skew-symmetric case, one identification was given by [FT19, Rem. 4.6]. Here, we slightly generalize it to the skew-symmetrizable case, and give another identification.
Definition 10.1.
For any cluster-positive element $M\in\mathcal{M}^{+}$, we define the class (A) and (B) by $\mathrm{Sk}(M)$ satisfies (A) and (B), respectively.
Lemma 10.2.
If a cluster-positive element $M\in M^{+}$ satisfies (B), the following statements hold.
(a) There exists a unique infinite sequence
(69) $M=M_{0}\overset{\gamma_{t_{1}}}{\mapsto}M_{1}\overset{\gamma_{t_{2}}}{\mapsto}M_{2}\overset{\gamma_{t_{3}}}{\mapsto}\cdots$
such that $M_{i-1}\nleq M_{i}$ for any $i=1,2,\dots$.
(b) This sequence has the limit $L(M)=\lim_{i\to\infty}M_{i}$. Moreover, $L(M)$ satisfies $\mathrm{Sk}(L(M))=(2,2,2)$. In particular, this limit $L(M)$ is a fixed point under the transformation $\Gamma$.
Proof.
Let $S=\mathrm{Sk}(M)$. Consider the sequence (47). Then, $S_{i-1}\nleq S_{i}$ holds. Set $M_{0}=M$ and $M_{i}=\gamma_{t_{i}}(M_{i-1})$. Then, by Lemma 9.4, this sequence also satisfies $M_{i-1}\nleq M_{i}$ as we desired. The uniqueness also follows from Lemma 9.4. (If there are another decreasing sequence, $S=\mathrm{Sk}(M)$ also has another decreasing sequence.) Thus, (a) holds. Next, we show (b). Let $M_{i}=\left(\begin{smallmatrix}a_{1;i}&a_{2;i}&a_{3;i}\\ a^{\prime}_{1;i}&a^{\prime}_{2;i}&a^{\prime}_{3;i}\end{smallmatrix}\right)$. By the assumption, each $a_{j;i}$ and $a^{\prime}_{j;i}$ ($j=1,2,3$) are monotonically decreasing for $i=0,1,2,\dots$. Moreover, since $M$ is cluster-positive, we have $a_{j;i},a^{\prime}_{j;i}\geq 0$. Thus, each component $a_{j;i},a^{\prime}_{j;i}$ should converge to some $u_{i},u^{\prime}_{i}\geq 0$, and we have $L(M)=\left(\begin{matrix}u_{1}&u_{2}&u_{3}\\ u^{\prime}_{1}&u^{\prime}_{2}&u^{\prime}_{3}\end{matrix}\right)$. Moreover, since $\mathrm{Sk}(S_{i})\overset{i\to\infty}{\longrightarrow}(2,2,2)$ by Lemma 7.7, we have $\sqrt{a_{j;i}a^{\prime}_{j;i}}\overset{i\to\infty}{\longrightarrow}2$. This implies $\sqrt{u_{j}u^{\prime}_{j}}=2$ ($j=1,2,3$) and $\mathrm{Sk}(L(M))=(2,2,2)$, which tells us that $L(M)$ is a fixed point by Lemma 9.8. ∎
By considering Lemma 7.7 and Lemma 8.2, we have the following theorem.
Theorem 10.3.
Let $M\in\mathcal{M}^{+}$ be cluster-positive element.
(a) This element $M$ satisfies (A) if and only if $C(M)<4$ or $\mathrm{Sk}(M)$ belongs to $\Gamma(p,p,2)$, $\Gamma(p,2,p)$, or $\Gamma(2,p,p)$ for some $p\geq 2$.
(b) This element $M$ satisfies (B) if and only if $C(M)=4$ and $\mathrm{Sk}(M)$ does not belong to $\Gamma(p,p,2)$, $\Gamma(p,2,p)$, and $\Gamma(2,p,p)$ for any $p\geq 2$. Moreover, in this case, there exists a decreasing sequence (69) uniquely, and it converges to some fixed point $L(M)$.
Proof.
Let $S=\mathrm{Sk}(M)$. By Lemma 7.7, if a cluster-positive element $S$ satisfies (B), then we have $C(S)=4$. Moreover, cluster-positive elements $S\in\mathcal{S}^{+}$ satisfying $C(S)=4$ and (A) should belong to $\Gamma(p,p,2)$, $\Gamma(p,2,p)$, or $\Gamma(2,p,p)$ by Lemma 8.2 (2). Thus, the claim holds. ∎